OpenStudy (anonymous):

amount of sales tax on a purchase varies directly as the price of the item varies. If a computer costs $400 before tax and the tax is $30 tax, which of the following equations is a correct step in solving for how much tax is needed for a computer that costs $500?

OpenStudy (anonymous):

400/30 = 500/y 400/30 = x/500 30/400 =500/y 30x = 200,000 These are my options. I know it's not the second or last one.

OpenStudy (solomonzelman):

If the comp costed 400 w/o tax, and with tax 430, find how much percent of the price is the tax. \[\frac{30}{400} \times 100--->\frac{15}{2}--->7.5\] so if the tax is 7.5% how much is 7.5% from 500?